# Leveraging Multi-Modal Information to Enhance Dataset Distillation

### v1

@@ 1 Introduction
Recent advances in dataset distillation have primarily focused on two key directions: matching-based methods [30, 27, 29, 1] and generative model priors [2, 17], both of which aim to generate synthetic datasets that closely approximate real data distributions. Matching-based methods optimize the alignment between real and synthetic images, ensuring that the distilled dataset preserves essential structural and statistical properties. These approaches have shown effectiveness in image-only datasets by directly optimizing synthetic samples against real data distributions. Meanwhile, generative model priors leverage GAN-based or diffusion models to synthesize high-quality samples, capturing complex variations in the data. While both techniques have demonstrated success in image-based tasks, they largely overlook the potential benefits of multi-modal information, such as textual descriptions and object-centric features, which could further enhance model generalization and feature learning.
To address these limitations, we propose a multi-modal dataset distillation framework that integrates diverse data modalities, such as caption descriptions, segmentation masks, and bounding boxes, into the distillation process. By incorporating multi-modal cues, our approach enables the model to learn from a more comprehensive representation of the dataset, rather than relying solely on image information. Caption features provide high-level semantic understanding, masks and bounding boxes help localize important regions within images, all contributing to a more informative synthetic dataset. This integration not only enhances feature learning but also improves model robustness in downstream tasks, making dataset distillation more scalable and effective beyond conventional image-based methods.
@@ 2 Related work
Dataset distillation was initially introduced together with model selection [21, 18]. Subsequently, matching methods have been explored that reduce the distance between the training process on synthetic images and the same teacher model trained on the real data, such as Dataset Condensation with Gradient Matching (DC) [30, 27, 25], Distribution Matching (DM) [29, 31], Matching Training Trajectories (MTT) [1], Sequential Subset Matching [5], and feature alignment of convolutional networks [20, 14]. To improve performance, other approaches, such as factorization [10], accumulated trajectory errors [4], calibration techniques [32], and Frequency Domain utilization [16] are also proposed. Because of the high-frequency noise in pixel space, [28, 2] focus on synthesizing images in the latent space using pre-trained GAN-type generative models. Other methods address different phases of dataset distillation, such as the concept of the distillation space [9], the clustering process that selects real images for the following matching [11], and new matching metrics with mutual information [15]. Sun et al. [19] proposed cropping patches from real images and concatenating high-scoring patches to generate synthetic images. Recently, several approaches have been proposed that involve training a diffusion model as part of the dataset distillation process. Su et al. [17] trained a latent diffusion model on prototypes rather than using traditional matching-based optimization, enabling the generation of more diverse and high-quality synthetic datasets. Gu et al. [6] fine-tuned a diffusion model to generate realistic images while introducing a minimax criterion to simultaneously improve representativeness and diversity, ensuring a more balanced and informative synthetic dataset. For large-scale datasets like the full ImageNet-1K, While existing methods [24, 23] have proposed learning image-text similarity as an auxiliary objective during dataset distillation for multi-modal data, we are, to the best of our knowledge, the first to rely on multi-modal data for distillation of the popular ImageNet-1K dataset, and demonstrate the superiority of our approach.
@@ 3.2 Caption combination
Figure 3: Overview of the Mask Matching Framework. (a) Masked Distribution Matching: The background is removed, retaining only the target object, which is then processed by a ConvNet. Distribution matching is applied to align the feature representations of real and synthetic images. (b) Masked Gradient Matching: gradient matching is applied, ensuring that the synthetic dataset preserves the underlying learning dynamics of the real dataset.
@@ 4.2 Cross architecture results
To assess the generalization ability of our dataset distillation approach, we conduct cross-architecture evaluations. The distillation phase is trained on a ConvNet, while the classification phase is evaluated on four different architectures: ResNet18, VGG11, ViT, and AlexNet. We report the average classification accuracy across these models to measure the transferability of our synthetic dataset across diverse network structures. Our experiments demonstrate that incorporating multi-modal information, including captions and masks, significantly enhances the quality of the distilled dataset and leads to improved performance in downstream classification tasks.

### v2

@@ 1 Introduction
Recent advances in dataset distillation have primarily focused on two key directions: matching-based methods [30, 27, 29, 1] and generative model priors [2, 17], both of which aim to generate synthetic datasets that closely approximate real data distributions. Matching-based methods optimize the alignment between real and synthetic images, ensuring that the distilled dataset preserves essential structural and statistical properties. These approaches have shown effectiveness in image-only datasets by directly optimizing synthetic samples against real data distributions. Meanwhile, generative model priors leverage GAN-based or diffusion models to synthesize high-quality samples, capturing complex variations in the data. While both techniques have demonstrated success in image-based tasks, they largely overlook the potential benefits of multi-modal information. Recent multi-modal methods [24, 23] have introduced image-text similarity objectives in vision language tasks. To the best of our knowledge, our approach is the first to comprehensively leverage multi-modal data, such as caption descriptions and object-centric features, from the widely-used ImageNet-1K dataset to assist the distillation process.
Building on this insight, we propose a multi-modal dataset distillation framework that integrates diverse data modalities, such as caption descriptions, segmentation masks, and bounding boxes, into the distillation process. By incorporating multi-modal cues, our approach enables the model to learn from a more comprehensive representation of the dataset, rather than relying solely on image information. Caption features provide high-level semantic understanding, masks and bounding boxes help localize important regions within images, all contributing to a more informative synthetic dataset. This integration not only enhances feature learning but also improves model robustness in downstream tasks, making dataset distillation more scalable and effective beyond conventional image-based methods.
@@ 2 Related work
Dataset distillation was initially introduced together with model selection [21, 18]. Subsequently, matching methods have been explored that reduce the distance between the training process on synthetic images and the same teacher model trained on the real data, such as Dataset Condensation with Gradient Matching (DC) [30, 27, 25], Distribution Matching (DM) [29, 31], Matching Training Trajectories (MTT) [1], Sequential Subset Matching [5], and feature alignment of convolutional networks [20, 14]. To improve performance, other approaches, such as factorization [10], accumulated trajectory errors [4], calibration techniques [32], and Frequency Domain utilization [16] are also proposed. Because of the high-frequency noise in pixel space, [28, 2] focus on synthesizing images in the latent space using pre-trained GAN-type generative models. Other methods address different phases of dataset distillation, such as the concept of the distillation space [9], the clustering process that selects real images for the following matching [11], and new matching metrics with mutual information [15]. Sun et al. [19] proposed cropping patches from real images and concatenating high-scoring patches to generate synthetic images. Recently, several approaches have been proposed that involve training a diffusion model as part of the dataset distillation process. Su et al. [17] trained a latent diffusion model on prototypes rather than using traditional matching-based optimization, enabling the generation of more diverse and high-quality synthetic datasets. Gu et al. [6] fine-tuned a diffusion model to generate realistic images while introducing a minimax criterion to simultaneously improve representativeness and diversity, ensuring a more balanced and informative synthetic dataset. For large-scale datasets like the full ImageNet-1K, While prior works have explored vision language methods [24, 23] on other datasets for dataset distillation. Our approach is the first to comprehensively leverage multi-modal data of the widely-used ImageNet-1K specifically for the dataset distillation process, demonstrating notable improvements over existing methods.
@@ 3.2 Caption combination
Figure 3: Overview of the Mask Matching Framework. (a) Masked Distribution Matching: Distribution matching is applied to align the feature representations of real and synthetic images. (b) Masked Gradient Matching: gradient matching is applied, ensuring that the synthetic dataset preserves the underlying learning dynamics of the real dataset.
@@ 4.2 Cross architecture results
To assess the generalization ability of our dataset distillation approach, we conduct cross-architecture evaluations. The distillation phase is trained on a ConvNet, while the classification phase is evaluated on four different architectures: ResNet18, VGG11, ViT, and AlexNet. We report the average classification accuracy across these models to measure the transferability of our synthetic dataset across diverse network structures.